OpenStudy (anonymous):

What polynomial has roots of -4, 1, and 6

OpenStudy (anonymous):

(x+1)(x+2)(x-3)=x3-7x-6

OpenStudy (anonymous):

@makaylarrrrrr

OpenStudy (anonymous):

A x3 - 3x2 - 22x + 24 B x3 - x2 - 26x - 24 C x3 + x2 - 26x + 24 D x3 + 3x2 + 14x - 24 @lupita1995

OpenStudy (anonymous):

what

OpenStudy (anonymous):

Be a LOT more careful. Not sure how I talked myself into that. Once we have the FACTORS from the ROOTS, THEN it's as obvious as I said. Root -4 1 6 Factor x - (-4) = x+4 x - 1 x - 6 Now, (+4)(-1)(-6) = +24 and immediately discard B and D.

OpenStudy (anonymous):

asically, the fact that -4, 1, and 6 are roots of this polynomial tell us that: (x−(−4))(x−1)(x−6) is the polynomial we are after. Your job is to multiply this all together. That will give you the answer.

OpenStudy (anonymous):

Note this is the same as: (x+4)(x−1)(x−6)

OpenStudy (anonymous):

im still not sure how to do that

OpenStudy (anonymous):

Be a little careful. In your example, the last term is (-a)(-b)(-c)(-d), and the negatives cancel out. Will that happen in our problem?